OpenStudy (anonymous):

Write the expression ln+ 3x^2/(x+3)^5 in an equivalent form that has no powers, products, or quotients in it. I got ln3x^2−ln(x+3)^5=2ln3x−5ln(x+3)

OpenStudy (anonymous):

"−5ln(x+3). The underlined section is correct. The other part is almost. You wrote it as if (3x)2 not as written 3x2. also break down the 3 x" I don't understand what I did wrong.

OpenStudy (e.mccormick):

3 is not squared in the first one.

OpenStudy (e.mccormick):

\[\frac{3x^2}{(x+3)^5}\implies \\ ln\left(\frac{3x^2}{(x+3)^5}\right)\implies \\ ln(3x^2)-ln(x+3)^5\implies \\ ln(3)+ln(x^2)-5 ln(x+3)\implies \\ ln3+2lnx-5 ln(x+3)\]

OpenStudy (e.mccormick):

It was not \((3x)^2\). So the 3 was not squared, it was multiplying.